OpenStudy (anonymous):

Use De Moivre's Theorem to find all of the complex solutions of the equation x³-i=0. Express the answer in rectangular form.

OpenStudy (anonymous):

havent i already done that?

OpenStudy (anonymous):

hi I wasn't sure about the e

OpenStudy (anonymous):

wt doubt about it?

OpenStudy (anonymous):

I didn't know how to ask you about it without re-posting...

OpenStudy (anonymous):

solving \[x^3=i\] yes?

OpenStudy (anonymous):

its done earlier satellite....he just has a doubt about the presentation

OpenStudy (anonymous):

I learnt it a different way and ended up with

OpenStudy (anonymous):

ooh sorry. rectangular means \[a+bi\] yes?

OpenStudy (anonymous):

(√3/2 + i1/2), -√3/2 + i1/2), (0-i) .... is this correct?

OpenStudy (anonymous):

I'm not sure

OpenStudy (anonymous):

angle for i is \[\frac{\pi}{2}\]so one cubed root will have angle \[\frac{\pi}{6}\]

OpenStudy (anonymous):

no..i gave u the euler forms of the three solns..can u retype em here for me plz..

OpenStudy (anonymous):

x = e^(ipi/6), e^(5ipi/6) and e^(7ipi/6)

OpenStudy (anonymous):

now can u convert all these into rectangular forms for me?

OpenStudy (anonymous):

\[cos(\frac{\pi}{6})+isin(\frac{\pi}{6})=\frac{\sqrt{3}}{2}+\frac{1}{2}i\]\]

OpenStudy (anonymous):

no that's what I was having trouble with

OpenStudy (anonymous):

the conversion

OpenStudy (anonymous):

for 5pi/6 x = cos(5pi/6) + isin(5pi/6)

OpenStudy (anonymous):

similarly for 7pi/6 x = cos(7pi/6) + isin(7pi/6)

OpenStudy (anonymous):

can u do that nw?

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

I really have no idea

OpenStudy (anonymous):

wts cos(5pi/6)? and sin 5pi/6?

OpenStudy (anonymous):

-√3/2 & 1/2

OpenStudy (anonymous):

righto....so x = -√3/2 + i(1/2) isnt it?

OpenStudy (anonymous):

right that's one of the answers I got

OpenStudy (anonymous):

now try for pi/6 in the same way

OpenStudy (anonymous):

along with (√3/2 + i1/2) & (0-i)

OpenStudy (anonymous):

how the hell did u get those two? try the same way for pi/6 and 7pi/6

OpenStudy (anonymous):

cos(pi/6) is √3/2 and sin(pi/6) is 1/2 right?

OpenStudy (anonymous):

right

OpenStudy (anonymous):

so x = √3/2 +i(1/2)

OpenStudy (anonymous):

?

OpenStudy (anonymous):

ys